OpenStudy (anonymous):

mily was born, she weighed 8.2 pounds. Two months later she weighed 10.15 pounds. How much weight did Emily gain in two months? A. 1.05 pounds B. 1.95 pounds C. 2.05 pounds D. 18.35 pounds

OpenStudy (anonymous):

1.95

OpenStudy (meehan98):

You subtract 10.15 from the initial weight (8.2 pounds) to get how much she gained.

OpenStudy (anonymous):

Which number sentence should be used to solve this problem? Paul used 15.55 feet of rope to make a tree swing. He used 27.42 feet of rope to make a rope ladder. How much rope did Paul use to make the tree swing and the rope ladder? A. 15.55 + 27.42 = ? B. 27.42 - 15.55 = ? C. 15.55 - 27.42 = ?

OpenStudy (meehan98):

If you're adding the feet of rope for the swing (15.55) + the feet of rope for the ladder (27.42), your number sentence would be: 15.55 + 27.42 =